the perimeter of a rectangle is 42m . the length of the rectangle is 3m less than twice the width . fine the length and the width of the rectangle .
Answer by actuary(112)

Let l = the length of the rectangle
Let w = the width of the rectangle
2*l+2*w=42 (Equation #1)
l = 2*w-3 (Equation #2)
Substitute equation 1 for "l" in Equation #2
2*(2w-3)+2w=42
4*w-6+2*w=42
Collect like terms
6w=48
Solve for w
w=8
Substitute w=8 into Equation #1
2*l+2*8=42
2*l+16=42
Solve for l
2*l=26
So l=13.